OpenStudy (anonymous):

A solution of salt and water contains 75 grams of water per 150 milliliters of the solution. If 1 mole of water weighs 16 grams, _____ moles of water would be present in 30 milliliters of the solution. [Round off your answer to two decimal places]

OpenStudy (anonymous):

i got .94

OpenStudy (anonymous):

\[\frac{ 75 \cancel{\text{ g H}_2 \text{O}} }{ 150 \cancel{\text{ ml sol'n} }}\cdot \frac{ 1 \text{ mol of H}_2 \text{O}}{ 16 \cancel{\text{ g H}_2 \text{O} }}\cdot 30 \cancel{\text{ ml sol'n}}\]

OpenStudy (anonymous):

=30/32 = .9375 mol = .94 mol. good job!

OpenStudy (anonymous):

can you help with one more

OpenStudy (anonymous):

Which table best represents the steps to solve the equation 8 + x over 3 = 4? Step Justification Step 1: x over 3 = 4 + 8 The same number +8 can be added to both sides of the equation Step 2: x over 3 = 12 Adding 8 and 4 on the right hand side Step 3: x = 12 • 3 The same number 3 can be multiplied on both sides of the equation Step 4: x = 36 Multiplying 12 and 3 on the right hand side Step Justification Step 1: x over 3 = 4 - 8 The same number -8 can be added to both sides of the equation Step 2: x over 3 = -4 Adding -8 and 4 on the right hand side Step 3: x = -4 • 3 The same number 3 can be multiplied on both sides of the equation Step 4: x = -12 Multiplying -4 and 3 on the right hand side Step Justification Step 1: x over 3 = 4 - 8 The same number - 8 can be added to both sides of the equation Step 2: x over 3 = -4 Adding 4 and -8 on the right hand side tep 3: x = - 4 - 3 The same number 3 can be subtracted from both sides of the equation Step 4: x = -7 Simplifying -4 -3 Step Justification Step 1: x over 3 = -4 - 8 The same number -8 can be added to both sides of the equation Step 2: x over 3 = -4 Adding -8 and 4 on the right hand side Step 3: x = -4 ÷ 3 The same number 3 can be divided on both sides of the equation Step 4: x = -1.33 Dividing -4 by 3 on the right hand side

OpenStudy (anonymous):

which do you think it is?

OpenStudy (anonymous):

i don't know.

OpenStudy (anonymous):

write out the equation, solve it, check it and then see which matches. that's what i did.

OpenStudy (anonymous):

\(8+\frac{x}{3}=4\)

OpenStudy (anonymous):

okay i don't know how, that is why i need help.

OpenStudy (anonymous):

to solve this, what's the first step?

OpenStudy (anonymous):

subtract ?

OpenStudy (anonymous):

subtract what?

OpenStudy (anonymous):

the 4 and 8

OpenStudy (anonymous):

and why?

OpenStudy (anonymous):

because whatever you do to one side you have to do the other.

OpenStudy (anonymous):

yes, but the goal here is to isolate the term with the x in it. as such, you subtract 8 from both sides. does that make sense?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so what is the new equation?

OpenStudy (anonymous):

x=-4

OpenStudy (anonymous):

no, \(\frac{x}{3}=-4\)

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

the i divide?

OpenStudy (anonymous):

then *

OpenStudy (anonymous):

no you multiply.

OpenStudy (anonymous):

and get 12